.a block of wood of 1.5kg mass floats on water with 68 percent of its volume submerged. a lead block is placed on the wood and the wood is the fully submerged with the lead entirely out of the water. find the mass of the lead block?

1
Expert's answer
2021-12-16T11:34:18-0500

M=1.5 kg

Vin = 68%

"M = ρ_{water} V_{in} \\

V_{in} = \frac{1.5}{1000} =1.5 \times 10^{-3} \;m^3 \\

V_o = \frac{V_{in}}{0.68} \\

V_o = 2.206 \times 10^{-3} \; m^3"

When the additional mass of lead is placed on the top, the remaining 32% of volume also goes into the water.

"m_{block}=1000 \times 0.32 \times 2.206 \times 10^{-3} = 0.706 \; kg"